Find a polynomial with roots of 5,-8 and 2 such that f(4)=-9

Thanks for your help

Answer by funmath(2933)

Find a polynomial with roots of 5,-8 and 2 such that f(4)=-9
:
y=a(x-b)(x-c)(x-d) where b,c, and d are the roots.
y=a(x-5)(x+8)(x-2) Now use f(4)=-9 to find a.
-9=a(4-5)(4+8)(4-2)
-9=a(-1)(12)(2)
-9=a(-24)
-9/-24=a(-24)/-24
9/24=a
:
Your polynomial in question is:
y=(9/24)(x-5)(x+8)(x-2) Most books and teachers are happy with this, but you could multiply it out if you want.
Happy Calculating!!!